A committee of five members is to be randomly selected from a group of nine freshman and seven sophomores. How many different committees of three freshmen and two sophomores can be chosen?
Answer by stanbon(75887)

A committee of five members is to be randomly selected from a group of nine freshman and seven sophomores. How many different committees of three freshmen and two sophomores can be chosen?
-----
# of 3 freshman groups = 9C3 = (9*8*7)/(1*2*3) = 3*4*7 = 84
# of 2 sophomore groups = 7C2 = (7*6)/(1*2) = 21
----
Ans:: 84*21 = 1764
-----------
Cheers,
Stan H.
-------